Threaded metric fasteners are available in several systems, two of which are given in Table 31.1: coarse and fine. Fine threads are not cut as deep and, as a result, have a larger tensile stress area for the same nominal diameter. Using a proof strength of 228 MPa and a 12 mm bolt, determine the (a) maximum preload that can be used for coarse pitch and fine pitch threads and (b) percent increase in preload of fine threads compared to course threads.
Answer #1

amate se pitah di =d-1-22 68 6 9 r -12-1.2268 a(S dp = d-c-644514f Caludak -tte dlmmtin プ the filth me aun A(10857) aleulab t2-0uas (1s) 11. 1881mm ไม่ pit, dianutn and mean mind dianeん) 4 X1oo